OpenStudy (anonymous):

find the intercepts of 4x+y=8

OpenStudy (goldphenoix):

x and y intercepts or just y?

OpenStudy (anonymous):

both

OpenStudy (goldphenoix):

Alright. To find y intercept, x must be 0. (0, y). Therefore, you must multiply x by 0. What does that mean? 4(0) + y = 8 What does that give you?

OpenStudy (goldphenoix):

To make it easier, what does 4 multiply by 0 give you?

OpenStudy (anonymous):

0

OpenStudy (goldphenoix):

Great. Now you have 7 = 8. In other words, the y-intercept is (0,8) not 8, since 8 is not a coordinate. ^_^ Now we basically do the same thing to find the x-intercept: 4x+(0)=8 What do you get?

OpenStudy (goldphenoix):

Opps. Now you have y = 8* My bad. >_>

OpenStudy (goldphenoix):

Anyways, since it's zero and is by itself, you can pretend that it doesn't exist. So now you have 4x = 8 We want 1x not 4x, so we must divide both side by 4. 8 divided by 4 = ?